# Delving into Temperature Scaling for Adaptive Conformal Prediction

### v2

@@ Abstract
[…]
@@ 1 Introduction
Ensuring the reliability of model predictions is crucial for the safe deployment of machine learning such as autonomous driving (Bojarski et al., 2016) and medical diagnostics (Caruana et al., 2015). Numerous methods have been developed to estimate uncertainty and incorporate it into predictive models, including confidence calibration (Guo et al., 2017) and Bayesian neural networks (Smith, 2013). However, these approaches lack theoretical guarantees of model performance. Conformal prediction, on the other hand, offers a systematic approach to construct prediction sets that contain ground-truth labels with a desired coverage guarantee (Vovk et al., 2005; Shafer & Vovk, 2008; Balasubramanian et al., 2014; Angelopoulos & Bates, 2021; Manokhin, 2022). This framework thus provides trustworthiness in real-world scenarios where wrong predictions are dangerous and costly.
@@
In the literature, conformal prediction is frequently associated with confidence calibration, which expects the model to predict softmax probabilities that faithfully estimate the true correctness (Wang et al., 2021; Wei et al., 2022; Yuksekgonul et al., 2023; Wang, 2023; Wang et al., 2024). For example, existing conformal prediction methods usually employ temperature scaling (Guo et al., 2017), a post-hoc method that rescales the logits with a scalar temperature, to process the model output for a better calibration performance (Angelopoulos et al., 2021b; Lu et al., 2022; 2023; Gibbs et al., 2023). The underlying hypothesis is that well-calibrated models could yield precise probability estimates, thus enhancing the reliability of generated prediction sets. However, the rigorous impacts of current confidence calibration techniques on conformal prediction remain ambiguous in the literature, which motivates our analysis of the connection between conformal prediction and confidence calibration.
@@
In this paper, we empirically show that existing methods of confidence calibration increase the size of prediction sets generated by adaptive conformal prediction methods. Moreover, high-confident predictions, rescaled by a small temperature value (Guo et al., 2017), often result in efficient prediction sets, while maintaining the desired coverage. To explain this phenomenon, we theoretically prove that a prediction, applied with a smaller temperature, could result in a more efficient prediction set on expectation. However, simply adopting an extremely small temperature value may result in invalid and meaningless prediction sets since some tail probabilities would be truncated to zero due to the finite precision problem. Given these findings, our goal is to automatically search for a temperature value that can improve the efficiency of prediction sets for adaptive conformal prediction methods.
To this end, we propose a variant of temperature scaling, Conformal Temperature Scaling (ConfTS), which optimizes the temperature value by minimizing the efficiency gap, i.e., the deviation between the threshold and the non-conformity score of the ground truth. We calculate the efficiency gap with the non-randomized APS score (Romano et al., 2020) for a hold-out dataset. In effect, ConfTS optimizes the temperature value to improve the efficiency of prediction sets, preserving the marginal coverage. Notably, our method is compatible with the original temperature scaling designed for confidence calibration, as we can pick temperature values according to the purpose during inference.
Extensive experiments show that ConfTS can effectively enhance existing adaptive conformal prediction techniques. In particular, our method drastically improves the efficiency of the prediction sets for APS (Romano et al., 2020) and RAPS (Angelopoulos et al., 2021b). For instance, using ViT-B-16 (Dosovitskiy et al., 2021) on ImageNet (Deng et al., 2009), ConfTS reduces the average set size of APS at $\alpha=0.1$ from $14.6$ to $2.3$, and declines that of RAPS from 6.9 to 1.8. Furthermore, ConfTS improves the conditional coverage of APS, and enhances the performance of the training-time method, ConfTr (Stutz et al., 2022). In practice, our approach is straightforward to implement within deep learning frameworks, requiring no additional computational costs on temperature scaling.
@@ 2 Preliminary
In this paper, we focus on adaptive conformal prediction methods, which are designed to guarantee conditional coverage by improving adaptiveness (Romano et al., 2020). However, they usually suffer from inefficiency in practice: these methods commonly produce large prediction sets (Angelopoulos et al., 2021b). In particular, we take the two representative methods: Adaptive Prediction Sets (APS) (Romano et al., 2020) and Regularized Adaptive Prediction Sets (RAPS) (Angelopoulos et al., 2021b).
@@
Notably, both methods incorporate a uniform random variable $u$ to achieve exact $1-\alpha$ coverage (Angelopoulos et al., 2021b). Moreover, we use coverage, average size, and size-stratified coverage violation (SSCV) (Angelopoulos et al., 2021b) to assess the marginal and conditional coverage, as well as the efficiency of prediction sets. A detailed description of the metrics is provided in Appendix A.
Table 1: The performance of APS and RAPS on CIFAR-100 and ImageNet dataset using various post-hoc calibration methods. In particular, we apply vector scaling (VS), Platt scaling (PS), and temperature scaling (TS). We do not employ calibration techniques in the baseline (Base). We repeat each experiment for 20 times. $\downarrow$” indicates smaller values are better. “” and “” indicate whether the performance is superior/inferior to the baseline. Bold numbers are superior results. The results show that all post-hoc confidence calibration methods deteriorate the efficiency of APS and RAPS.
Datasets Metrics ResNet18 ResNet50 ResNet101
Base VS PS TS Base VS PS TS Base VS PS TS
CIFAR-100 Accuracy 0.76 0.75 0.76 0.76 0.77 0.77 0.77 0.77 0.78 0.79 0.78 0.78
ECE(%) $\downarrow$ 5.68 3.67 4.20 4.29 8.79 3.62 3.81 4.06 10.8 3.27 3.83 3.62
APS Coverage 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
Average size $\downarrow$ 8.73 8.94 10.1 10.0 4.91 6.69 7.75 7.35 4.01 5.77 6.99 6.66
RAPS Coverage 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
Average size $\downarrow$ 4.14 4.40 4.71 4.67 2.63 3.58 3.72 3.85 2.27 3.39 3.54 3.53
ImageNet Accuracy 0.69 0.68 0.69 0.69 0.76 0.75 0.76 0.76 0.77 0.76 0.77 0.77
ECE(%) $\downarrow$ 2.63 2.14 2.10 2.27 3.69 1.50 2.24 2.35 5.08 1.38 2.02 2.20
APS Coverage 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
Average size $\downarrow$ 14.1 17.3 15.9 16.0 9.06 12.0 12.0 12.1 6.95 11.1 10.7 10.6
RAPS Coverage 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90 0.90
Average size $\downarrow$ 9.61 10.6 11.75 11.30 5.99 7.56 7.52 7.16 4.82 6.86 6.85 6.59
@@ 3.1 Adaptive conformal prediction with calibrated prediction
Confidence calibration methods deteriorate the efficiency of prediction sets.
To figure out the correlation between confidence calibration and adaptive conformal prediction, we incorporate various confidence calibration methods to adaptive conformal predictors for classification models on CIFAR-100 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009). Specifically, we use 6 calibration methods, including 4 post-hoc methods – vector scaling (Guo et al., 2017), Platt scaling (Platt et al., 1999), temperature scaling (Guo et al., 2017), Bayesian methods (Daxberger et al., 2021), and 2 training methods – label smoothing (Szegedy et al., 2016), mixup (Zhang et al., 2018). More details of calibration methods and setups are presented in Appendix B and Appendix C, respectively.
[…]
@@ 3.2 Adaptive Conformal prediction with high-confidence prediction
High confidence enhances the efficiency of adaptive conformal prediction.
In Figures 1(a) and 1(b), we present the average size of prediction sets generated by APS and RAPS under various temperature values $t$. The results show that a highly-confident model, produced by a small temperature value, would decrease the average size of prediction sets. For example, using VGG16, the average size is reduced by four times – from 20 to 5, with the decrease of the temperature value from 1.3 to 0.5. There naturally arises a question: is it always better for efficiency to take smaller temperature values?
@@
In Figure 1(c), we report the average size of prediction sets produced by APS on ImageNet with ResNet18, using extremely small temperatures (i.e. $t\in\{0.12,0.14,\cdots,0.2\}$). Different from the above, APS generates larger prediction sets with smaller temperatures in this range, even leading to conservative coverage. This problem stems from floating point numerical errors caused by finite precision (see Appendix E for a detailed explanation). The phenomenon indicates that it is non-trivial to find the optimal temperature value for the highest efficiency of adaptive conformal prediction.
@@ 3.3 Theoretical explanation
For instance $\bm{x}\in\mathcal{X}$, let $\mathcal{S}(\bm{x},k,t)$ be the non-conformity score function of an arbitrary class $k\in\mathcal{Y}$, defined as in Eq. 7. Then, for a fixed temperature $t_{0}$ and $\forall t\in(0,t_{0})$, we have
@@
The proof is provided in Appendix F.1. In Proposition 3.1, we show that the APS score increases as temperature decreases, and vice versa. Then, for any data sample $\bm{x}$ and a fixed temperature $t_{0}$, we further define $\epsilon(k,t)=\mathcal{S}(\bm{x},k,t)-\mathcal{S}(\bm{x},k,t_{0})\geq 0$ as the difference of the APS scores. As a corollary of Proposition 3.1, we conclude that $\epsilon(k,t)$ is negatively correlated with the temperature $t$. We provide the proof for this corollary in Appendix F.2. The corollary is formally stated as follows:
@@
In the following, we further explore how the change in the APS score affects the average size of the prediction set. In the theorem, we make two continuity assumptions on the CDF of the non-conformity score (see Appendix F.3), following prior works (Lei, 2014; Sadinle et al., 2019). Given these assumptions, we can derive an upper bound for the expected size of $\mathcal{C}(\bm{x},t)$ for any $t\in(0,t_{0})$:
@@
Under assumptions in Appendix F.3, there exists constants $c_{1},\gamma\in(0,1]$ such that
@@
Interpretation. The proof of Theorem 3.3 is presented in Appendix F.3. Through Theorem 3.3, we show that for any temperature $t$, the expected size of the prediction set $\mathcal{C}(\bm{x},t)$ has an upper bound with respect to the non-conformity score deviation $\epsilon$. Recalling that $\epsilon$ increases with the decrease of temperature $t$, we conclude that a lower temperature $t$ results in a larger difference $\epsilon$, thereby narrowing the prediction set $\mathcal{C}(\bm{x},t)$. Overall, the analysis shows that high-confidence predictions, produced by a small temperature, could lead to efficient prediction sets on expectation. Given the theoretical analysis, we propose to enhance the efficiency of adaptive conformal prediction by tuning the temperature. We proceed by introducing our method – Conformal Temperature Scaling.
4 Method
In the previous analysis, we show that temperature scaling optimized by negative log-likelihood deteriorates the efficiency of adaptive conformal prediction, while a small temperature can improve. Nevertheless, results in Figure 1(c) demonstrate that finding the optimal temperature for the highest efficiency is a non-trivial task. In this work, we propose Conformal Temperature Scaling, a variant of temperature scaling, to select an appropriate temperature for enhancing adaptive conformal prediction.
@@ 4 Method
Specifically, the optimal prediction set is the smallest set that allows the inclusion of the ground-truth label. The concept of optimal prediction set naturally leads to a way for quantifying the redundancy of generated prediction set $\mathcal{C}(\bm{x})$: we can compute the deviation between the size of the prediction set and that of the optimal prediction set $|\mathcal{C}(\bm{x})|-|\mathcal{C}^{*}(\bm{x})|$. However, it is challenging to perform optimization with the size difference due to its discrete property. To circumvent the issue, we convert the optimization objective of our method into a continuous loss function that is end-to-end trainable.
Efficiency gap. Recall that the prediction set is established through the $\tau$ calculated from the calibration set (Eq. (3)), the optimal set can be attained if the threshold $\tau$ well approximates the non-conformity score of the ground-truth label $\mathcal{S}(\bm{x},y)$. Therefore, we can also measure the redundancy of the prediction set by the differences between thresholds $\tau$ and the score of true labels, defined as:
@@
where $\mathcal{D}_{\mathrm{loss}}$ denotes the subset for computing ConfTS loss. Trained with the ConfTS loss, we can optimize the temperature $t$ for adaptive prediction sets with high efficiency without violating coverage. Since the threshold $\tau$ is determined by the pre-defined $\alpha$, our ConfTS method can yield different temperature values for each $\alpha$. Notably, ConfTS offers compelling advantages as a post-hoc method:
  • Algorithm-agnostic: ConfTS trained with APS score can improve other adaptive conformal prediction methods (e.g., RAPS) and is compatible with training-time methods such as ConfTr (Stutz et al., 2022) for improved efficiency. This is supported by Table 2 and Table 3.
  • Easy-to-use: Our method enhances the efficiency of conformal prediction in a parameter-efficient fashion. This stands in contrast to training methods (Stutz et al., 2022), which require optimizing the full parameters of networks and may degrade the accuracy. Moreover, our ConfTS is free from hyper-parameter tuning, and requires low computational resources.
  • Flexible: Our method does not conflict with confidence calibration, as it only replaces the temperature value. During inference, one may use different temperature values according to the objective, whether for improved calibration performance or efficient prediction sets.

### v3

# Does confidence calibration improve conformal prediction?

@@ Abstract
Conformal prediction is an emerging technique for uncertainty quantification that constructs prediction sets guaranteed to contain the true label with a predefined probability. Previous works often employ temperature scaling to calibrate classifiers, assuming that confidence calibration benefits conformal prediction. However, the specific impact of confidence calibration on conformal prediction remains underexplored. In this work, we make two key discoveries about the impact of confidence calibration methods on adaptive conformal prediction. Firstly, we empirically show that current confidence calibration methods (e.g., temperature scaling) typically lead to larger prediction sets in adaptive conformal prediction. Secondly, by investigating the role of temperature value, we observe that high-confidence predictions can enhance the efficiency of adaptive conformal prediction. Theoretically, we prove that predictions with higher confidence result in smaller prediction sets on expectation. This finding implies that the rescaling parameters in these calibration methods, when optimized with cross-entropy loss, might counteract the goal of generating efficient prediction sets. To address this issue, we propose Conformal Temperature Scaling (ConfTS), a variant of temperature scaling with a novel loss function designed to enhance the efficiency of prediction sets. This approach can be extended to optimize the parameters of other post-hoc methods of confidence calibration. Extensive experiments demonstrate that our method improves existing adaptive conformal prediction methods in classification tasks, especially with LLMs.
@@ 1 Introduction
Ensuring the reliability of model predictions is crucial for the safe deployment of machine learning such as autonomous driving (Bojarski et al., 2016) and medical diagnostics (Caruana et al., 2015). Numerous methods have been developed to estimate uncertainty and incorporate it into predictive models, including confidence calibration (Guo et al., 2017) and Bayesian neural networks (Smith, 2013). However, these approaches lack theoretical guarantees of model performance. Conformal prediction, on the other hand, offers a systematic approach to construct prediction sets that contain ground-truth labels with a desired coverage guarantee (Vovk et al., 2005; Shafer & Vovk, 2008; Balasubramanian et al., 2014; Angelopoulos & Bates, 2021). This framework thus provides trustworthiness in real-world scenarios where wrong predictions are dangerous.
@@
In the literature, conformal prediction is frequently associated with confidence calibration, which expects the model to predict softmax probabilities that faithfully estimate the true correctness (Wei et al., 2022; Yuksekgonul et al., 2023; Wang, 2023; Wang et al., 2024). For example, existing conformal prediction methods usually employ temperature scaling (Guo et al., 2017), a post-hoc method that rescales the logits with a scalar temperature, for a better calibration performance (Angelopoulos et al., 2021b; Lu et al., 2022; 2023; Gibbs et al., 2023). The underlying hypothesis is that well-calibrated models could yield precise probability estimates, thus enhancing the reliability of generated prediction sets. However, the rigorous impacts of current confidence calibration techniques on conformal prediction remain ambiguous in the literature, which motivates our analysis of the connection between conformal prediction and confidence calibration.
@@
In this paper, we empirically show that existing methods of confidence calibration increase the size of prediction sets generated by adaptive conformal prediction methods. Moreover, we find that predictions with high confidence (rescaled with a small temperature value) tend to produce efficient prediction sets while maintaining the desired coverage guarantees. However, simply adopting an extremely small temperature value may result in meaningless prediction sets, as some tail probabilities can be truncated to zero due to the finite-precision issue. Theoretically, we prove that a smaller temperature value leads to larger non-conformity scores, resulting in more efficient prediction sets on expectation. This highlights that rescaling parameters of post-hoc calibration methods, optimized by the cross-entropy loss, might counteract the goal of generating efficient prediction sets.
To validate our theoretical findings, we propose a variant of temperature scaling, Conformal Temperature Scaling (ConfTS), which rectifies the optimization objective through the efficiency gap, i.e., the deviation between the threshold and the non-conformity score of the ground truth. In particular, ConfTS optimizes the temperature value by minimizing the efficiency gap. This approach can be extended to optimize the parameters of other post-hoc methods of confidence calibration, e.g., Vector Scaling and Platt Scaling. Extensive experiments show that ConfTS can effectively enhance the efficiency of existing adaptive conformal prediction techniques, APS (Romano et al., 2020) and RAPS (Angelopoulos et al., 2021b). Notably, we empirically show that post-hoc calibration methods optimized by our loss function can also improve the efficiency of prediction sets in both image and text classification (including large language models), which demonstrates the generality of our method. In addition, we provide an ablation study of loss functions to show that the proposed loss function can outperform the ConfTr loss (Stutz et al., 2022). In practice, our approach is straightforward to implement within deep learning frameworks, requiring no hyperparameter tuning and additional computational costs compared to standard temperature scaling.
We summarize our contributions as follows:
  • We discover that current confidence calibration methods typically lead to larger prediction sets in adaptive conformal prediction, while high-confidence predictions (using small temperatures) can enhance the efficiency of prediction sets. We further identify a practical limitation where extremely small temperature values cause numerical precision issues.
  • We provide a theoretical analysis by proving that applying smaller temperature values in temperature scaling results in more efficient prediction sets on expectation. This theoretical insight explains the relationship between confidence calibration and conformal prediction.
  • We validate our theoretical findings by developing Conformal Temperature Scaling (ConfTS), a variant of temperature scaling that exploits the relationship between temperature and set efficiency. Extensive experiments demonstrate that ConfTS enhances the efficiency of prediction sets in adaptive conformal prediction and can be extended to other post-hoc methods of confidence calibration.
@@ 2 Preliminary
In this paper, we focus on adaptive conformal prediction methods, which are designed to improve conditional coverage (Romano et al., 2020). However, they usually suffer from inefficiency in practice: these methods commonly produce large prediction sets (Angelopoulos et al., 2021b). In particular, we take the two representative methods: APS (Romano et al., 2020) and RAPS (Angelopoulos et al., 2021b).
@@
Notably, both methods incorporate a uniform random variable $u$ to achieve exact $1-\alpha$ coverage (Angelopoulos et al., 2021b). Moreover, we use coverage and average size to evaluate the prediction sets. A detailed description of the metrics is provided in Appendix B.
@@ 3.1 Adaptive conformal prediction with calibrated prediction
To figure out the correlation between confidence calibration and adaptive conformal prediction, we incorporate various confidence calibration methods to adaptive conformal predictors for a ResNet50 model (He et al., 2016) on CIFAR-100 dataset (Krizhevsky et al., 2009). Specifically, we use six calibration methods, including four post-hoc methods – vector scaling (Guo et al., 2017), Platt scaling (Platt et al., 1999), temperature scaling (Guo et al., 2017), Bayesian methods (Daxberger et al., 2021), and two training methods – label smoothing (Szegedy et al., 2016), mixup (Zhang et al., 2018). More details of calibration methods and setups are presented in Appendix C and Appendix D.
Table 1: The performance of APS and RAPS on CIFAR-100 dataset with ResNet50 model, using various calibration methods. In particular, we apply label smoothing (LS), Mixup (Mixup), Bayesian methods (Bayesian), vector scaling (VS), Platt scaling (PS), and temperature scaling (TS). We do not employ calibration techniques in the baseline (Base). We repeat each experiment for 20 times. $\downarrow$” indicates smaller values are better. “” and “” indicate whether the performance is superior/inferior to the baseline. The results show that existing confidence calibration methods deteriorate the efficiency of APS and RAPS.
Method Base LS Mixup Bayesian TS PS VS
Accuracy 0.77 0.78 0.78 0.77 0.77 0.77 0.77
ECE $\downarrow$ 8.79 4.39 2.96 4.30 3.62 3.81 4.06
$\alpha=0.1$ APS Coverage 0.90 0.90 0.90 0.90 0.90 0.90 0.90
Avg.size $\downarrow$ 4.91 11.9 12.5 7.55 6.69 7.75 7.35
RAPS Coverage 0.90 0.90 0.90 0.90 0.90 0.90 0.90
Avg.size $\downarrow$ 2.56 9.50 10.2 6.46 3.58 3.72 3.85
$\alpha=0.05$ APS Coverage 0.95 0.95 0.95 0.95 0.95 0.95 0.95
Avg.size $\downarrow$ 11.1 19.8 20.1 15.6 12.8 13.9 11.3
RAPS Coverage 0.95 0.95 0.95 0.95 0.95 0.95 0.95
Avg.size $\downarrow$ 6.95 14.5 15.5 9.34 10.4 11.0 8.70
Confidence calibration methods deteriorate the efficiency of prediction sets. In Table 1, we present the performance of confidence calibration and conformal prediction using APS and RAPS with various calibration methods for a ResNet50 model. The results show that the influences of those calibration methods are consistent: models calibrated by both post-hoc and training calibration techniques generate large prediction sets with lower ECE (i.e., better calibration). For example, on the ImageNet dataset, temperature scaling enlarges the average size of prediction sets of APS from 9.06 to 12.1, while decreasing the ECE from 3.69% to 2.24%. This finding demonstrates an inverse relationship between calibration performance and prediction set efficiency. In addition, incorporating calibration methods into conformal prediction does not violate the $1-\alpha$ marginal coverage as the assumption of data exchangeability is still satisfied: we use a hold-out validation dataset for conducting confidence calibration methods. In addition, we present the results of the LAC score (Sadinle et al., 2019) in Appendix E.1, where we observe no clear correlation between confidence calibration methods and conformal prediction.
@@ 3.2 Adaptive Conformal prediction with high-confidence prediction
High confidence enhances the efficiency of adaptive conformal prediction. In Figures 1(a) and 1(b), we present the average size of prediction sets generated by APS and RAPS under various temperature values $t$. The results show that a highly-confident model, produced by a small temperature value, would decrease the average size of prediction sets. For example, using VGG16, the average size is reduced by four times – from 20 to 5, with the decrease of the temperature value from 1.3 to 0.5. There naturally arises a question: is it always better for efficiency to take smaller temperature values?
@@
In Figure 1(c), we report the average size of prediction sets produced by APS on ImageNet with ResNet18, using extremely small temperatures (i.e. $t\in\{0.12,0.14,\cdots,0.2\}$). Different from the above, APS generates larger prediction sets with smaller temperatures in this range, even leading to conservative coverage. This problem stems from floating point numerical errors caused by finite precision (see Appendix F for a detailed explanation). The phenomenon indicates that it is non-trivial to find the optimal temperature value for the highest efficiency of adaptive conformal prediction.
@@ 3.3 Theoretical explanation
For instance $\bm{x}\in\mathcal{X}$, let $\mathcal{S}(\bm{x},k,t)$ be the non-conformity score function of an arbitrary class $k\in\mathcal{Y}$, defined as in Eq. 7. Then, for a fixed temperature $t_{0}\in\mathbb{R}^{+}$ and $\forall t\in(0,t_{0})$, we have
@@
The proof is provided in Appendix G.1. In Proposition 3.1, we show that the APS score increases as temperature decreases, and vice versa. Then, for a fixed temperature $t_{0}\in\mathbb{R}^{+}$, we further define $\epsilon(k,t)=\mathcal{S}(\bm{x},k,t)-\mathcal{S}(\bm{x},k,t_{0})\geq 0$ as the difference of the APS scores. As a corollary of Proposition 3.1, we conclude that $\epsilon(k,t)$ is negatively correlated with the temperature $t$. We provide the proof for this corollary in Appendix G.2. The corollary is formally stated as follows:
@@
In the following, we further explore how the change in the APS score affects the average size of the prediction set. In the theorem, we make two continuity assumptions on the CDF of the non-conformity score (see Appendix G.3), following prior works (Lei, 2014; Sadinle et al., 2019). Given these assumptions, we can derive an upper bound for the expected size of $\mathcal{C}(\bm{x},t)$ for any $t\in(0,t_{0})$:
@@
Under assumptions in Appendix G.3, there exists constants $c_{1},\gamma\in(0,1]$ such that
@@
Interpretation. The proof of Theorem 3.3 is presented in Appendix G.3. Through Theorem 3.3, we show that for any temperature $t$, the expected size of the prediction set $\mathcal{C}(\bm{x},t)$ has an upper bound with respect to the non-conformity score deviation $\epsilon$. Recalling that $\epsilon$ increases with the decrease of temperature $t$, we conclude that a lower temperature $t$ results in a larger difference $\epsilon$, thereby narrowing the prediction set $\mathcal{C}(\bm{x},t)$. Overall, the analysis shows that tuning temperature values can potentially enhance the efficiency of adaptive conformal prediction. In practice, we may employ grid search to find the optimal $T$ for conformal prediction, but it requires defining the search range of $T$ and cannot be extended to post-hoc calibration methods with more parameters, like Platt scaling and Vector scaling. Thus, we propose an alternative solution for automatically optimizing the parameters to enhance the efficiency of conformal prediction.
3.4 An alternative method for improving efficiency
In the previous analysis, we empirically and theoretically demonstrate that standard temperature scaling optimized by negative log-likelihood often leads to degraded efficiency, while searching for a relatively small temperature can potentially address this issue. In this work, we propose an alternative method, Conformal Temperature Scaling (ConfTS), to automatically optimize the parameters of post-hoc calibration methods. This is a variant of temperature scaling that directly optimizes the objective function toward generating efficient prediction sets, and it can be extended to other post-hoc calibration methods.
@@ 3.4 An alternative method for improving efficiency
Specifically, the optimal prediction set is the smallest set that allows the inclusion of the ground-truth label. Recall that the prediction set is established through the $\tau$ calculated from the calibration set (Eq. (3)), the optimal set can be attained if the threshold $\tau$ well approximates the non-conformity score of the ground-truth label $\mathcal{S}(\bm{x},y)$. Therefore, we can measure the redundancy of the prediction set by the differences between thresholds $\tau$ and the score of true labels, defined as:
@@
where $\mathcal{D}_{\mathrm{loss}}$ denotes the subset for computing ConfTS loss. Trained with the ConfTS loss, we can optimize the temperature $t$ for adaptive prediction sets with high efficiency without violating coverage. Since the threshold $\tau$ is determined by the pre-defined $\alpha$, our ConfTS method can yield different temperature values for each $\alpha$.
Extensions to other post-hoc calibration methods. Noteworthy, our ConfTS loss is a general method and can be easily incorporated into existing post-hoc calibration methods such as Platt scaling (Platt et al., 1999) and vector scaling (Guo et al., 2017). Formally, for any rescaling function $\phi_{\bm{\theta}}$ with parameters $\bm{\theta}$, we can formulate the method as follows. First, we define the $k$-th softmax probability after rescaling as:
$\pi_{k}(\bm{x};\bm{\theta})=\sigma(\phi_{\bm{\theta}}\cdot f(\bm{x}))_{k}=\frac{e^{[\phi_{\bm{\theta}}\cdot f(\bm{x})]_{k}}}{\sum_{i=1}^{K}e^{[\phi_{\bm{\theta}}\cdot f(\bm{x})]_{i}}}$
The corresponding non-conformity score for each class $k\in\mathcal{Y}$ is given by $\mathcal{S}(\bm{x},k;\bm{\theta})=\sum_{i=1}^{k}\pi_{k}(\bm{x};\bm{\theta})$. With the threshold $\tau(\bm{\theta})$, we rewrite the ConfTS loss by
$\mathcal{L}_{\mathrm{ConfTS}}(\bm{x},y;t)=(\tau(\bm{\theta})-\mathcal{S}(\bm{x},y,\bm{\theta}))^{2},$
Then, the optimization objective can be formulated as:
$\bm{\theta^{*}}=\underset{\bm{\theta}}{\arg\min}\frac{1}{|\mathcal{D}_{\mathrm{loss}}|}\sum_{(\bm{x}_{i},y_{i})\in\mathcal{D}_{\mathrm{loss}}}\mathcal{L}_{\mathrm{ConfTS}}(\bm{x}_{i},y_{i};\bm{\theta}).$